OpenStudy (anonymous):

A hamster ball has a radius of 15 centimeters. What is its volume written in terms of π? 3,375π cm3 300π cm3 4,500π cm3 2,531π cm3

OpenStudy (kropot72):

The formula for the volume of a sphere is:\[(3\div4)\times \pi \times r ^{3}\] Substituting for r = 15: \[(15)^{3}\times (3\div4)\times \pi\] \[2,531\pi cm ^{3}\]

OpenStudy (anonymous):

ty :D

OpenStudy (kropot72):

You're welcome :)